A company is planning a $80 million expansion. The expansion is to be financed by selling $30 million in new debt and $50 million in new common stock. The before-tax required rate of return on debt is 8 percent and the required rate of return on equity is 16 percent. If the company is in the 40 percent tax bracket, what is the firm's cost of capital??
Answer #1

After tax cost of debt=8(1-0.4)=4.8%

Hence firm's cost of capital=(Respective costs*Respective investment weights)

=(30/80*0.048)+(50/80*0.16)

=11.8%